### v3

# Analytically Continuing the Randomized Measurement Toolbox

Abstract
We develop a framework for extracting non-polynomial analytic functions of density matrices in randomized measurement experiments by a method of analytical continuation. A central advantage of this approach, dubbed stabilized analytic continuation (SAC), is its robustness to statistical noise arising from finite repetitions of a quantum experiment, making it well-suited to realistic quantum hardware. As a demonstration, we use SAC to estimate the von Neumann entanglement entropy of a numerically simulated quenched Néel state from Rényi entropies estimated via the randomized measurement protocol. We then apply the method to experimental Rényi data from a trapped-ion quantum simulator, extracting subsystem von Neumann entropies at different evolution times. Finally, we briefly note that the SAC framework is readily generalizable to obtain other nonlinear diagnostics, such as the logarithmic negativity and Rényi relative entropies.
Introduction.– Recent advances in quantum simulator platforms enable the study of complex quantum many-body phenomena, such as thermalization, scrambling, and topological order, in well-controlled laboratory settings [44, 2]. Understanding these phenomena requires access to the entanglement structure of quantum states. Quantum information theory provides a quantitative framework for characterizing entanglement through the von Neumann entanglement entropy of a density matrix $\rho$,
$S(\rho)=-\mathrm{Tr}(\rho\log_{2}\rho),$ (1)
which serves as a fundamental measure of bipartite entanglement [25]. It plays a central role in diagnosing topological order [31, 33], probing thermalization and information scrambling [49, 19, 13, 30, 38], and studying quantum criticality [53, 40] and the emergence of bulk gravity in AdS/CFT [48, 47]. However, directly measuring the von Neumann entropy for large quantum systems in quantum simulators remains experimentally challenging, motivating the search for scalable measurement approaches.
While full quantum state tomography allows complete reconstruction of a quantum state and thereby access to its von Neumann entropy, its sample complexity scales sharply exponentially with system size, rendering it impractical beyond few-qubit systems [22, 39]. Recent advances in randomized measurement (RM) protocols [17, 26, 14] have pushed the boundary to substantially larger system sizes for a broad class of polynomial functions of the density matrix [54]. Such functions can also be estimated utilizing many-body interference between multiple physical copies of quantum state prepared simultaneously in a quantum experiment [27, 29]. These protocols enable measurement of scalable quantities such as Rényi entropies $S_{k}(\rho)=\frac{1}{1-k}\log_{2}\mathrm{Tr}\rho^{k}$, $k\geq 2$ integer [16, 17], Hilbert–Schmidt fidelities, and mixed-state entanglement witnesses [15]. Such quantities serve as powerful diagnostics of correlations and entanglement [5, 15], but they are not entanglement measures [25]. Moreover, Rényi entropies can display qualitatively different behavior from the von Neumann entropy [45, 55, 4, 1]. Therefore, reliable estimates of the von Neumann entropy are essential for a quantitative characterization of entanglement in many-body quantum systems.
Figure 1: SAC and traditional fitting methods [52, 35] estimating the half-chain von Neumann entropy in the ground state of the transverse-field Ising model $H=-J\sum_{i}\sigma_{i}^{z}\sigma_{i+1}^{z}-h\sum_{i}\sigma_{i}^{x}$ away from criticality ($J=1$, $h=0.5$, 15 sites). Error percentage of different methods are shown as a function of the largest Rényi index in the dataset. Left: for noiseless Rényi inputs. Right: for inputs with 10% independent Gaussian noise, averaged over 200 realizations.
In this Letter, we develop a framework to reliably estimate the von Neumann entropy in quantum simulation experiments using a finite number of measured Rényi entropies. While the von Neumann entropy can, in principle, be obtained from knowledge of all integer Rényi entropies via analytic continuation, $S_{\mathrm{vN}}=\lim_{k\to 1^{+}}S_{k}$, a consequence of Carlson’s theorem [23], in practice only a limited number of Rényi orders are accessible, each with finite statistical uncertainty. This renders the continuation problem ill-posed and highly sensitive to statistical noise—much like the analytic continuation of Green’s functions away from the Matsubara domain or extracting physical scattering amplitudes in quantum chromodynamics [21, 28, 50, 12, 42]. We employ the stabilized analytic continuation (SAC) approach of Ciulli and Spearman [8, 9, 6, 7, 20], which reformulates analytic continuation as a constrained convex optimization problem: among all analytic functions compatible with the data, SAC selects the one minimizing a chosen $L^{2}$-norm. Our novel adaption of SAC yields stable and noise-resilient estimates of the von Neumann entropy even when only a few Rényi entropies are available. Fig. 1 shows the performance of SAC in an example of estimating the half-chain von Neumann entropy of a spin chain model, and compares it against more traditional extrapolation schemes. For benchmarking, we synthetically add independent Gaussian noise to each input Rényi entropies. Beyond entanglement entropy, our framework applies broadly to estimating general non-polynomial spectral functions of quantum states accessible through analytic continuation.
In the following, we present our adaptation of the SAC method, including deriving an explicit analytic expression of the von Neumann entropy estimate from a few Rényi entropies, and generalizing to the case with (correlated) statistical noise. We benchmark our framework using numerically simulated Rényi data for a quenched Néel state and demonstrate its performance on experimental data from a trapped-ion quantum simulator [5], where it enables extraction of subsystem von Neumann entropies across time.
Stabilized analytic continuation.– Let $S_{z}(\rho)=\frac{1}{1-z}\log_{2}\mathrm{Tr}\rho^{z}$, $z\in\mathbb{C}$, denote the Rényi function on the complex plane. Suppose we are provided noisy estimates of this function at $N$ integer points $z=2,3,.\cdots,k_{\text{max}}$. Our task is to analytically continue this dataset to the von Neumann point, $z=1$. Assuming the true underlying function is smooth, traditional approaches typically involve fitting with lower-degree polynomials to mitigate overfitting [24, 36]. Nevertheless, choosing a low-degree polynomial still requires model assumptions and entails an arbitrary selection of polynomial basis.
This leads us to the method of stabilized analytic continuation (SAC), which bypasses these challenges by exploiting the fundamental analytic properties of the underlying function [8, 9, 6, 7, 20]. To adapt it to the von Neumann entropy problem, we construct a variational function whose minimum $L^{2}$-norm determines the optimal estimate of the von Neumann entropy. We explicitly derive the analytical expression for the von Neumann estimate resulting from this variational approach in the noiseless case. We also extend the SAC framework, originally formulated for uncorrelated noise, to handle correlated noise. Our adaptation is found to be more robust to realistic noise than conventional extrapolation techniques, as confirmed using data with correlated noise below.
In the first part of the analysis, we assume the data points are noiseless. We begin by defining the following discrepancy function on the $z$ plane
$D_{\alpha}(z)=\frac{S_{z}(\rho)}{z-1}-\frac{\alpha}{z-1}$ (2)
where $\alpha$ is a real variational parameter. The discrepancy function assumes the values $\{d_{i}(\alpha)=\frac{S_{i}(\rho)}{z_{i}-1}-\frac{\alpha}{z_{i}-1}\}$ at the data points $\{z_{i}=2,\cdots,k_{\text{max}}\}$. Note that upon dividing $S_{z}(\rho)$ by (z1)(z-1), we have artificially introduced a simple pole at the von Neumann point, $z=1$, thereby enhancing the structure of the function in the vicinity of this point. The residue of $D_{\alpha}(z)$ at this pole is $S_{vN}(\rho)-\alpha$. If $\alpha=S_{vN}(\rho)$, the pole cancels out altogether.
We now map the analytic domain of $D_{\alpha}(z)$ to the interior of the unit disk via a conformal transformation $w(z)$, chosen such that
  1. 1.
    $D_{\alpha}(z(w))$ remains analytic for $|w|<1$
  2. 2.
    The real half-line $1,\infty)\in\mathbb{R}_{z}$ is mapped to the interval $1,1]\in\mathbb{R}_{w}$
  3. 3.
    The point $z=1$ is mapped to w=1w=-1.
[…]
$\xi=\cosh(\frac{z-1}{\epsilon}+\frac{i\pi}{2}),\mbox{ followed by }\hskip 5.0ptw=\frac{\xi-\eta i}{\xi+\eta i}$ (3)
Here $\eta$ is a free parameter. In practice, $\eta$ and $\epsilon$ are chosen based on benchmarks on the simulation data, and these choices correspond to when the pole contribution dominates the norm in (4) below. The Rényi function within the unit disk $S_{w}\equiv S_{z(w)}$ now assumes the same values $\{S_{2}(\rho),S_{3}(\rho),\cdots,S_{k_{\text{max}}}(\rho)\}$ at the new points $w_{i}=w(i)$ for $i=2,\cdots,k_{\text{max}}$. Since $S_{w}$ is free of singularities when $|w|<1$, so is $D_{\alpha}(w)\equiv D_{\alpha}(z(w))$. However, $D_{\alpha}(w)$ still has a pole on the unit circle at the new von Neumann point, w=1w=-1. We can “measure” the amount of structure of the function on the unit circle by defining a suitable norm. An convenient choice is the $L^{2}$ (pseudo-) norm  [^1] We remark here that the $L^{2}$ (pseudo-)norm defined in (4) is not the only choice one can make. In fact, there are four different $L^{2}$ type norms one can write down where the integrand is expressed as the square of the real part of the function, the imaginary part of the function, the angular derivative of the real part of the function or the angular derivative of the imaginary part of the function. The choice (4) is motivated by the observation that the pole at w=1w=-1 only manifests as a divergence in the imaginary part of the function.:
[…]
Note that G=0||G||=0 does not imply $G=0$; it only implies $G(w)$ must be a constant in the interior of the disk. We can restrict to the set of functions which vanish at some point $w_{0}$ in the disk, then (4) defines a valid norm. Often, we choose $w_{0}=w_{2}$, the location of the first input data point, and replace $D_{\alpha}(w)$ with $D^{\prime}_{\alpha}(w)=D_{\alpha}(w)-d_{2}(\alpha)$.
Crucially, if $\alpha\neq S_{vN}(\rho)$, $D^{\prime}_{\alpha}(w)||$ will be very large due to the presence of the singularity on the unit circle. We can therefore estimate the optimal $\alpha_{\min}=S_{vN}(\rho)$ by carrying out the following dual minimization steps:
  1. 1.
    For fixed $\alpha$, we search for the analytic function $Y_{\alpha}(w)$ which takes the desired values $\{Y_{\alpha}(w_{i})=D^{\prime}_{\alpha}(w_{i})=d_{i}(\alpha)-d_{2}(\alpha)\}$ at the points $\{w_{i}\}$ and minimizes the norm $Y_{\alpha}||$. This minimization can be recast into a linear matrix optimization problem by leveraging the analyticity of $Y_{\alpha}(w)$ [8]. The minimal norm
[…]
    provides a measure of the structure that is forced on the function by the data itself.
  2. 2.
    Next we minimize $\delta(\alpha)$ over all possible values of $\alpha$. This is where we expect the pole at w=1w=-1 cancels out, and the minimal value $\alpha_{\min}$ serves as our best estimate for $S_{vN}(\rho)$.
Due to the absence of noise, both of these minimization steps can be performed exactly to obtain the following closed form expression for the von Neumann estimate:
$S^{\text{est}}_{vN}=\frac{\sum_{i,j=3}^{k_{\text{max}}}(A^{-1})_{ij}\Bigl(\frac{S_{i}(\rho)}{i-1}-S_{2}(\rho)\Bigl)\Bigl(\frac{1}{j-1}-1\Bigl)}{\sum_{i,j=3}^{k_{\text{max}}}(A^{-1})_{ij}\Bigl(\frac{1}{i-1}-1\Bigl)\Bigl(\frac{1}{j-1}-1\Bigl)}$ (5)
where $A_{ij}$ is a symmetric and positive definite matrix (see Supplemental Material for its expression).
In the presence of noise, Step 1 above is modified to include sampling of all data points that lie within a suitable neighbourhood of the mean values $D^{\prime}_{\alpha}(w_{j})$. If $C$ denotes the covariance matrix of the data points, then the goodness of fit of an arbitrary point $\{y_{j}\}$ in data space can be measured using a $\chi^{2}$ statistic: $\chi^{2}(\mathbf{y};\alpha)=\sum_{i,j=2}^{k_{\text{max}}}(y_{i}-D^{\prime}_{\alpha}(w_{i}))(C^{\prime-1})_{ij}(y_{j}-D^{\prime}_{\alpha}(w_{j}))$, where $\mathbf{y}=\{y_{2},...,y_{k_{\text{max}}}\}$, and $C^{\prime}_{ij}=\frac{C_{ij}}{(z_{i}-1)(z_{j}-1)}$ is the covariance matrix of the corresponding discrepancy function. Note that for the noisy data case, we treat the subtraction point as a variational parameter as well [10, 11], and hence the sum runs from $i,j=2$. The minimal norm $\delta(\alpha)$ is now computed by minimizing over all possible points y subject to the constraint $\chi^{2}(\mathbf{y};\alpha)\leq\chi^{2}_{0}$ for some constant $\chi^{2}_{0}$ typically chosen to be $\mathcal{O}(k_{\mbox{\tiny max}})$. Namely,
[…]
Step 2 of the minimization procedure stays the same. In this case, we are unable to provide a closed form expression for $S^{\text{est}}_{vN}(\rho)$, but we can still efficiently estimate this value numerically for arbitrarily large systems (see Supplemental Material for more details).
Figure 2: Numerical benchmarks using simulated data for a 10-qubit Néel state quenched under Hamiltonian (8) with decoherence included. Left: von Neumann entropy estimates from different methods at fixed time, t=5t=5ms, and subsystem size, $L=5$ sites. Center: von Neumann entropy versus time for $L=5$. Right: vN entropy versus subsystem size at t=5t=5ms.
Randomized measurements and classical shadow tomography.— We briefly summarize the randomized measurement (RM) protocol [17, 26, 14]. For a system of $L$ qubits in an unknown state $\rho$, local random unitaries $U=U_{1}\otimes\cdots\otimes U_{L}$ are drawn independently from a unitary 3-design [14] and followed by projective measurements in the computational basis. From the resulting bit strings, collected over $N_{u}$ unitaries and $N_{m}$ repetitions, one constructs classical shadows $\rho^{(m)}$}, which are unbiased estimators of the density matrix, $\mathbb{E}[\rho^{(m)}]=\rho$. Expectation values of observables and polynomial functions of $\rho$, such as trace moments $p_{k}=\mathrm{Tr}(\rho^{k})$, can then be estimated via U-statistics [15]:
[…]
where $I_{k}$ denotes distinct index sets of size $k$. Because evaluating $\hat{p}_{k}$ scales as $\mathcal{O}(N_{u}^{k})$, implying significant post-processing time, the batch-shadow variant [46] groups $N_{u}$ shadows into $N_{B}\ll N_{u}$ batches, averaging within each batch before computing the U-statistic. This strongly reduces the computational cost while retaining near-optimal statistical accuracy [46]; we employ this variant throughout, typically with $N_{B}=10\sim 20$.
While in principle Rényi entropies up to order $N_{u}$ ($N_{B}$) could be estimated from the data, higher-order Rényi entropies suffer from rapidly increasing statistical uncertainty [15], yielding in practice a finite set of accessible Rényi orders. In addition, shadow tomography cannot directly estimate non-polynomial functionals such as the von Neumann entropy. In the following, we show that SAC allows accurate inference from finite-order Rényi entropies (up to $k_{\text{max}}=6$) obtained via shadow tomography, yielding robust von Neumann entropy estimates for both simulated and experimental data.
Benchmarking SAC with numerical simulations.— We demonstrate the performance of the SAC method on synthetic data generated from numerical simulations of quench dynamics in a 10-qubit system. The system is initialized in a low-entropy Néel state and subsequently evolved under a spin-$1/2$ Hamiltonian [5, 43],
$H=\sum_{i<j}J_{ij}(\sigma_{i}^{+}\sigma_{j}^{-}+\sigma_{i}^{-}\sigma_{j}^{+})+B\sum_{j}\sigma_{j}^{z},$ (8)
where $\sigma_{i}^{\pm}$ are spin ladder operators and the couplings follow an approximate power-law decay $J_{ij}\approx J/|i-j|^{1.2}$ with $J=420\mathrm{s}^{-1}\ll B$. The simulations incorporate realistic decoherence sources, including imperfect Néel-state preparation, spontaneous emission, spin flips, and depolarization from local unitaries [5].
RM are simulated on a classical computer to construct classical (batch) shadows, which are then used to estimate Rényi entropies [15]. We perform 1000 independent shadow experiments, each consisting of $N_{u}=500$ random unitaries and $N_{m}=150$ measurements per unitary. Using U-statistics, Rényi entropies of orders $k=2,\cdots,6$ are computed for each experiment. The von Neumann entropy is extracted with both SAC and traditional fitting schemes, including Chebyshev polynomial interpolation [35] and least-squares fitting [52]. These estimates are compared with the exact von Neumann entropy obtained directly from the simulation (Fig. 2). For SAC, the 1000 Rényi estimates are grouped into 50 groups of 20 Rényi sets each. For each group, we calculate the mean and covariance matrix of the 20 Rényi sets and feed these to the noisy SAC protocol. The resulting 50 von Neumann entropy estimates are then averaged to obtain the final mean and statistical uncertainty. Results across different subsystem sizes and timesteps are summarized in Fig. 2.
Application to Trapped-ion Quantum Simulator.– Finally, we apply SAC to extract the von Neumann entropy from Rényi data obtained in a trapped-ion experiment with ten $^{40}\text{Ca}^{+}$ ions, prepared in an approximate Néel state and evolved under Hamiltonian (8). The experimental dataset comprises a single shadow tomography experiment with $N_{u}=500$ and $N_{m}=150$ for various timesteps and subsystem sizes. Jackknife resampling is applied to obtain the mean Rényi entropies for orders $k=2,\cdots,6$ and the covariance matrix [^2] The mean and covariance matrix here are jackknife-corrected.. While, naively, the Jackknife resampling would require $\mathcal{O}(N_{B})$ evaluations of the U-statistic, we employ here a fast Jackknife implementation based on hashing. These values are used as input for the SAC protocol to produce the von Neumann entropy estimate shown in Fig. 3.
The challenge in this experimental application lies in estimating the error bar for the von Neumann entropy. Unlike the simulation with 1000 independent shadow experiments, our dataset is derived from only one. We explored several standard methods for error analysis, including bootstrapping the 500 density matrices and splitting them into smaller groups. However, both approaches proved unreliable: the bootstrapping produced unrealistically small error bars, while the splitting yielded results that were highly dependent on how the density matrices were grouped. Failure of these methods indicate a larger $N_{u}$ is required to obtain a robust estimate of the error bar. However, even with a moderate increase to $N_{u}=1000$, we can achieve a meaningful estimation: by splitting the data into two independent groups of 500 unitaries, the difference between the resulting von Neumann estimates provides an approximation of the true error bar.
Figure 3: (a)(b): $S_{vN}$ as a function of time $t$ and subsystem size $L$ obtained via SAC in a trapped-ion experiment with ten $^{40}\text{Ca}^{+}$ ions, initialized in an approximate Néel state and evolved under Hamiltonian (8). For comparison, $S_{vN}$ obtained via SAC using simulation data, as well as Rényi data from simulated randomized measurements, are plotted in (c)(d). In (a)(c), $L=5$. In (b)(d), t=5t=5ms.
Discussion.– We introduced a robust framework for extracting the von Neumann entropy from finite sets of Rényi entropies subject to statistical noise obtained via randomized measurements. When tested on extensive simulation data, our method outperforms conventional fitting approaches. We applied our method to experimental data from a trapped-ion quantum simulator, obtaining direct estimates of subsystem von Neumann entropy growth following a quantum quench. Since randomized measurements are now routinely used to measure entropies in many platforms, our noise-resilient approach provides a practical post-processing tool for analyzing data affected by statistical uncertainty and may substantially reduce measurement overheads. It also extends to situations where Renyi entropies are measured through many-body interference experiments [27, 29].
There are many potential applications of SAC beyond extracting von Neumann entropies. The method naturally extends to any information-theoretic quantity accessible through a replica trick—such as logarithmic negativity, Rényi relative entropies, Petz–Rényi divergences, or Uhlmann fidelity [15, 34, 51, 41, 37, 56, 3, 32]—which we leave for future exploration. More broadly, SAC offers a promising strategy for addressing long-standing analytic-continuation challenges in many other areas of physics, from quantum Monte Carlo studies to the reconstruction of scattering amplitudes.
Acknowledgement.– We thank Hsin-Yuan Huang, Nima Lashkari, Jong Yeon Lee, Eric Schultz, Vandana Ramakrishnan, and Daniel Mark for helpful discussions. JKF is supported by the Marvin L. Goldberger Member Fund at the Institute for Advanced Study and the National Science Foundation under Grant No. PHY-2514611. BV acknowledges support from the hybdrid quantum initiative (HQI ANR-22-PNCQ-0002).
Note: The open-source code to extract Rényi entropies from batch shadows, applied to the experimental data of Ref. [5], is available on Github, and presented in details in Ref. [18].
Appendix A Stabilized Analytic Continuation
In this appendix, we review the key features of the SAC method outlined in references [8, 9, 6, 7, 20]. Additional details can be found in the original references.
The basic setup of the problem is as described in the main text. $F(w)$ is an unknown analytic function that is holomorphic in the interior of the unit disk. It assumes real values on the real axis: $F(w_{i})=a_{i}\in\mathbb{R}$ for $w_{i}\in(-1,1)$ and i=2,3,,N+1i=2,3,...,N+1, where $N$ denotes the number of data points ($N=k_{\text{max}}-1$ in the main text). Our aim is to fit for the value of the function at a specific point on the unit circle, which we take to be w=1w=-1 without loss of generality. SAC was originally developed by Ciulli and Spearman [8, 9, 6, 7, 20] to diagnose resonances of the Green’s function which lie outside, but close to the unit disk, using a finite set of noisy data points provided within the interior of the disk. The presence of nearby poles results in a significant amount of structure in the function on the unit circle, which can be measured by a suitable choice of norm. Minimizing this norm over all consistent analytic continuations then yields the minimal amount of structure that is forced upon the function by the data itself.
To adapt this method to our target problem, we first artificially introduce a pole at w=1w=-1 by defining $G(w)=\frac{F(w)}{w+1}$. The aim now is to estimate the residue of the pole at w=1w=-1. To do so, as explained in the main text, we define the discrepancy function $D_{\alpha}(w)=G(w)-\frac{\alpha}{w+1}$ where $\alpha$ is a variational parameter. If $\alpha=F(-1)$, the residue of $D_{\alpha}(w)$ is 0 at w=1w=-1, hence $D_{\alpha}(w)$ becomes analytic even on the unit circle. To measure the structure of the function near w=1w=-1, we introduce a norm defined entirely in terms of the values of the function on the unit circle. Given either a Dirichlet or a Neumann boundary problem, there are two corresponding unique norm choices. Both are continuous $L^{2}$ type norms which stem from an inner product. For the problem of analytic continuation of Rényi entropies, a proper choice of the norm is
$||X||=\frac{1}{2\pi}\int_{0}^{2\pi}\bigg|\frac{\partial}{\partial\theta}\text{Im}X(e^{i\theta})\bigg|^{2}d\theta=\frac{1}{2\pi}\int_{0}^{2\pi}\bigg|\frac{\partial}{\partial r}\text{Re}X(re^{i\theta})\big|_{r=1}\bigg|^{2}d\theta$ (9)
where the latter equality follows from the Cauchy-Riemann equations. Importantly, this norm is singular when the function $X$ possesses a pole at w=1w=-1. For the case of a constant non-zero function $X$, $X||$ is zero and is a pseudo-norm. To rule out this special case, we restrict $X$ to be the subset of analytic functions with $X(w_{0})=0$, where $w_{0}$ is an arbitrary subtraction point. In what follows, we replace $D_{\alpha}(w)$ with $D^{\prime}_{\alpha}(w)=D_{\alpha}(w)-D_{\alpha}(w_{0})$. A convenient choice for $w_{0}$ is $w_{2}$, the location of the first data point. Let us introduce the notation
$x_{,r}(\theta)=\frac{\partial\text{Re}X(re^{i\theta})}{\partial r}\bigg|_{r=1}$ (10)
As illustrated in [8, 9, 6, 7, 20], knowing $x_{,r}(\theta)$ on the unit circle allows one to reconstruct the function $X(w)$ in the interior of the unit disk upto an overall constant through the formula
$X(w)=X(w^{\prime})-\frac{1}{\pi}\int_{0}^{2\pi}x_{,r}(\theta)\ln|\frac{e^{i\theta}-w}{e^{i\theta}+w^{\prime}}\Bigg|d\theta$ (11)
Furthermore, if $w^{\prime}$ is chosen to be subtraction point $w_{0}$, then $x_{,r}(\theta)$ fully determines $X(w)$ in the interior of the disk.
In our case, $X(w)=D^{\prime}_{\alpha}(w)$, the subtracted discrepancy function. $D^{\prime}_{\alpha}(w)$ is required to take on the values
$d^{\prime}_{\alpha;i}=\bigg(\frac{a_{i}}{w_{i}+1}-\frac{a_{1}}{w_{0}+1}\bigg)-\bigg(\frac{\alpha}{w_{i}+1}-\frac{\alpha}{w_{0}+1}\bigg)$ (12)
at the data points $\{w_{i}\}$. There is an infinite family of analytic functions $D^{\prime}_{\alpha}(w)$ that satisfy the condition. From this infinite family, we need to select the analytic continuation that has the smallest norm. As illustrated in [8, 9, 6, 7, 20], this non-linear minimization problem can be recast into a linear matrix optimization problem, and the minimal norm (squared) $\delta_{\min}^{2}$ of $D^{\prime}_{\alpha}(w)$ is given by
$\delta_{\min}^{2}=\sum_{i,j=3}^{N+1}(A^{-1})_{ij}d^{\prime}_{\alpha;i}d^{\prime}_{\alpha;j}$ (13)
Here the sum starts from $i,j=3$ because we have chosen the subtraction point to be the first data point $w_{0}=w_{2}$. $A$ is an $(N-1)\times(N-1)$ positive-definite, symmetric matrix with matrix elements defined as
$A_{ij}=\frac{2}{\pi}\int_{0}^{2\pi}\ln|\frac{e^{i\theta}-w_{i}}{e^{i\theta}-w_{2}}\Bigg|\ln|\frac{e^{i\theta}-w_{j}}{e^{i\theta}-w_{2}}\Bigg|d\theta$ (14)
Note $A_{ij}$ only depends on the data points (locations) $w_{i}$ and not the data values itself. $\delta_{\min}^{2}$ still depends on $\alpha$ and as we vary $\alpha$, it should have a sharp minimum when $\alpha$ equals the true residue of $G(w)$, namely $\alpha=F(-1)$. $\delta_{\min}^{2}$ turns out to be a quadratic function of $\alpha$, and we can immediately write down a closed form expression for $\alpha$ that minimizes this quadratic function:
$\alpha_{\min}=\frac{\sum_{i,j=3}^{N+1}(A^{-1})_{ij}\Bigl(\frac{a_{i}}{w_{i}+1}-\frac{a_{1}}{w_{2}+1}\Bigl)\Bigl(\frac{1}{w_{j}+1}-\frac{1}{w_{2}+1}\Bigl)}{\sum_{i,j=3}^{N+1}(A^{-1})_{ij}\Bigl(\frac{1}{w_{i}+1}-\frac{1}{w_{2}+1}\Bigl)\Bigl(\frac{1}{w_{j}+1}-\frac{1}{w_{2}+1}\Bigl)}$ (15)
This is our estimate for the target value F(1)F(-1).
Appendix B Stabilized Analytic Continuation with Errors
We now consider the case where the prescribed data values $\{d_{i}\equiv d^{\prime}_{\alpha;i}\}$ are noisy. The associated errors are generically correlated and can be described by a symmetric, positive semi-definite covariance matrix $C$:
$C=\sum_{I=1}^{N}\epsilon^{2}_{I}{\bf e}_{I}{\bf e}_{I}^{T}$ (16)
where ${\bf e}_{I}$ denote the orthonormal eigenvectors and $\epsilon^{2}_{I}\geq 0$ denote the non-negative eigenvalues. Equivalently, we may write $C=ODO^{T}$ where $D=\text{diag}(\epsilon^{2}_{1},\cdots,\epsilon^{2}_{N})$ and $O=({\bf e}_{1},\cdots,{\bf e}_{N})^{T}$
The $N$ data points are now viewed as a point in $\mathbb{R}^{N}$, denoted by $\mathbf{d}=(d_{1},d_{2},\cdots,d_{N})^{T}$. An arbitrary point in this data space is denoted as $\mathbf{y}=(y_{1},y_{2},\cdots,y_{N})^{T}$. The goodness of fit of an arbitrary data point with the specified data values is measured by a $\chi^{2}$ statistic:
$\chi^{2}=(\mathbf{y}-\mathbf{d})^{T}C^{-1}(\mathbf{y}-\mathbf{d})=\sum_{I=1}^{N}\frac{(y_{I}-d_{I})^{2}}{\epsilon_{I}^{2}}$ (17)
In the second equality, we have expanded both $\mathbf{d}=\sum_{I=1}^{N}d_{I}{\bf e}_{I}$ and $\mathbf{y}=\sum_{I=1}^{N}y_{I}{\bf e}_{I}$ in the eigenbasis ${\bf e}_{I}$ of the covariance matrix $C$. We shall restrict ourselves to the set of points which satisfy $\chi^{2}\leq\chi^{2}_{0}$ for some chosen constant $\chi^{2}_{0}$. These points constitute an ellipoid in data space whose principal axes correspond to ${\bf e}_{I}$.
Due to errors in the data points, it now no longer makes sense to choose the subtraction point to correspond to one of the data points. To incorporate errors, the subtraction point can instead be chosen as the guess function’s value at $w_{2}$ [9], or as an additional variational parameter [20]. Here we opt for the latter and take $w_{0}$, an arbitrary point on the real interval (1,1)(-1,1), as our subtraction point. Given any data point $\mathbf{y}$ within this ellipse $\chi^{2}\leq\chi^{2}_{0}$, the minimal (squared) norm $\delta^{2}_{\min}$ is given by
$\delta^{2}_{\min}=(\mathbf{y}-\mathbf{y}_{0})^{T}A^{-1}(\mathbf{y}-\mathbf{y}_{0})$ (18)
Here $\mathbf{y}_{0}=y_{0}\mathbf{1}$ where $\mathbf{1}=(1,1,\cdots,1)^{T}$ and $y_{0}$ is the value of the guess function at $w_{0}$. Same as $w_{0}$, $y_{0}$ is also a free parameter. Again it is convenient to express this quantity in the ${\bf e}_{I}$ basis. To this end, let us first expand the $\mathbf{1}$ vector as $\mathbf{1}=\sum_{I=1}^{N}1_{I}{\bf e}_{I}$ where $1_{I}={\bf e}_{I}^{T}\bf{1}$. Next we define a new matrix $B=O^{T}AO$. Due to the orthogonality of $O$, $B^{-1}=O^{T}A^{-1}O$. In components, $(B^{-1})_{IJ}={\bf e}_{I}^{T}A^{-1}{\bf e}_{J}$. This matrix is clearly symmetric and positive definite. Then, the minimal norm $\delta^{2}_{\min}$ can be written as
$\delta^{2}_{\min}=\sum_{IJ=1}^{N}(B^{-1})_{IJ}(y_{I}-{y}_{0}1_{I})(y_{J}-{y}_{0}1_{J})$ (19)
Now the goal is to find the smallest value of the norm within the domain $\chi^{2}\leq\chi^{2}_{0}$. Due to convexity of $\delta^{2}_{\min}$, the minimum value is attained on the boundary of this domain which satisfies $\chi^{2}=\chi^{2}_{0}$. Thus, we can recast the problem of finding the data point satisfying $\chi^{2}=\chi^{2}_{0}$ which minimizes $\delta^{2}_{\min}$ as a Lagrange multiplier problem.
$G[\mathbf{y},y_{0},\lambda]=\delta^{2}_{\min}[\mathbf{y},y_{0}]+\lambda(\chi^{2}[\mathbf{y}]-\chi^{2}_{0})$ (20)
We then have the following two equations: $\frac{\partial G}{\partial y_{0}}=\frac{\partial G}{\partial y_{I}}=0$. These equations simplify to
$\sum_{J=1}^{N}B^{-1}_{IJ}(y_{J}-y_{0}1_{J})+\frac{\lambda}{\epsilon_{I}^{2}}(y_{I}-d_{I})=0$ ${\implies}(y_{I}-y_{0}1_{I})+\lambda\sum_{J=1}^{N}B_{IJ}\frac{(y_{J}-d_{J})}{\epsilon_{J}^{2}}=0$ (21)
$\sum_{I,J=1}^{N}1_{I}B^{-1}_{IJ}(y_{J}-y_{0}1_{J})=0$ ${\implies}\sum_{I=1}^{N}\frac{1_{I}(y_{I}-d_{I})}{\epsilon_{I}^{2}}=0$ (22)
It is now convenient to define the following variables
$p_{J}=\frac{d_{J}-y_{J}}{\epsilon_{J}}\qquad,\qquad q_{J}=\frac{d_{J}-y_{0}1_{J}}{\epsilon_{J}}\qquad,\qquad M_{IJ}=\frac{B_{IJ}}{\epsilon_{I}\epsilon_{J}}$ (23)
The Lagrange equations now simplify to
$(q_{I}-p_{I})-\lambda\sum_{J=1}^{N}M_{IJ}p_{J}=0$ (24)
$\sum_{I=1}^{N}\frac{1_{I}p_{I}}{\epsilon_{I}}=0$ (25)
In terms of these variables, note that $\delta^{2}_{\min}=\sum_{IJ=1}^{N}M^{-1}_{IJ}(q_{I}-p_{I})(q_{J}-p_{J})=\lambda^{2}\sum_{I,J=1}^{N}M_{IJ}p_{I}p_{J}$ whereas $\chi^{2}=\sum_{I=1}^{N}p_{J}^{2}$.
Let $\{\sigma_{r}\}$ and $\{\textbf{f}_{r}\}$ denote the eigenvalues and eigenvectors of $M_{IJ}$. Then expanding p and q as follows $\textbf{p}=\sum_{r=1}^{N}p_{r}\textbf{f}_{r}$ and $\textbf{q}=\sum_{r=1}^{N}q_{r}\textbf{f}_{r}$ allows us to re-express the first Lagrange equation as $q_{r}-p_{r}-\lambda\sigma_{r}p_{r}=0$ or $p_{r}=\frac{q_{r}}{1+\lambda\sigma_{r}}$. All that remains is to solve for $y_{0}$. We do so by defining 2 new vectors m and n as $\textbf{m}=\sum_{I=1}^{N}\frac{d_{I}}{\epsilon_{I}}\textbf{e}_{I}=\sum_{r=1}^{N}m_{r}\textbf{f}_{r}$ and $\textbf{n}=\sum_{I=1}^{N}\frac{1_{I}}{\epsilon_{I}}\textbf{e}_{I}=\sum_{r=1}^{N}n_{r}\textbf{f}_{r}$. Then notice that $q_{r}=m_{r}-y_{0}n_{r}$ and $p_{r}=\frac{m_{r}}{1+\lambda\sigma_{r}}-y_{0}\frac{n_{r}}{1+\lambda\sigma_{r}}$. The second Lagrange equation then implies
$0=\sum_{r}n_{r}p_{r}=\frac{\sum_{r}n_{r}m_{r}}{1+\lambda\sigma_{r}}-y_{0}\frac{\sum_{r}n_{r}^{2}}{1+\lambda\sigma_{r}}$
$\implies y_{0}=\frac{\frac{\sum_{r}n_{r}m_{r}}{1+\lambda\sigma_{r}}}{\frac{\sum_{r}n_{r}^{2}}{1+\lambda\sigma_{r}}}$ (26)
The only undetermined parameter is the Lagrange multiplier $\lambda$. It is fixed by imposing the condition $\chi^{2}=\sum_{r=1}^{N}p_{r}^{2}(\lambda)=\chi^{2}_{0}$. This is the only step that needs to be performed numerically. Once $\lambda$ has been determined, we can compute the minimal norm as $\delta^{2}_{\min}=\lambda^{2}\sum_{r=1}^{N}\sigma_{r}p^{2}_{r}$. Now, just as in the noiseless case, $\delta^{2}_{\min}$ is really a function of $\alpha$ since the data values $d_{i}\equiv d^{\prime}_{i;\alpha}$ are functions for $\alpha$. Thus, we need to solve for the value of $\alpha$ numerically which minimizes $\delta^{2}_{\min}$. This provides an estimate of the residue F(1)F(-1) in the noisy case.